### v1

# Nonlocal van der Waals density functional made faster

Abstract
A simplification of the VV10 van der Waals density functional [J. Chem. Phys. 133, 244103 (2010)] is made by an approximation of the integrand of the six-dimentional integral in terms of a few products of three-dimensional density-like distributions and potential-like functions of the interelectronic distance only, opening the way for its straightforward computation by fast multipole methods. An even faster computational scheme for molecular systems is implemented where the density-like distributions are fitted by linear combinations of usual atom-centered basis functions of Gaussian type and the six-dimensional integral is then computed analytically, at a fraction of the overall cost of a typical calculation. The simplicity of the new approximation is commensurate with that of the original VV10 functional, and the same level of accuracy is seen in tests on molecules.
I Introduction
Density functional theory is where the worlds of mathematics, physics, and chemistry meet in the most fruitful way, the rigorous theorems are followed by models ranging from approximations and interpolations between known limits down to empirical parametrizations of all kinds, to come up with a computational tool that helps the chemist to understand the structure and behavior of molecular systems and even to design new ones. Starting from the uniform electron gas Hohenberg and Kohn (1964); Kohn and Sham (1965), the generalized-gradient approximations Langreth and Mehl (1983); Becke (1986); Perdew and Wang (1986); Perdew, Burke, and Ernzerhof (1996) have brought the accuracy to a chemically meaningful level; a mixing of exact exchange Becke (1993a, b) or an even better long-range exchange correction Iikura et al. (2001); Heyd, Scuseria, and Ernzerhof (2003); Henderson, Janesko, and Scuseria (2008); Laikov (2019a) is the next step in reaching near chemical accuracy. What is still missing is the small but sometimes important dispersion (also known as van der Waals) energy contribution for which density functionals evolved Rydberg et al. (2000, 2003); Dion et al. (2004) in the form of a six-dimensional integral — a line of further simplifications Vydrov, Wu, and Van Voorhis (2008); Vydrov and Van Voorhis (2009, 2010a) (notwithstanding formal controversies Langreth and Lundqvist (2010); Vydrov and Van Voorhis (2010b)) has led to an elegant analytical expression for its integrand — but even then its direct numerical evaluation is rather time-consuming. In plane-wave methods a fast Fourier transform techniques Román-Pérez and Soler (2009); Sabatini, Gorni, and de Gironcoli (2013) can help, but here we want to deal with isolated molecular systems without any further atoms-in-molecules Gräfenstein and Cremer (2009); Steinmann and Corminboeuf (2010) kind of approximation.
Here we report a new way we have found to simplify the evaluation of the six-dimensional integral of dispersion energy functionals by approximating the integrand in terms of a few products of density-like distributions and potential-like functions — making it computable by fast multipole methods Rokhlin (1985) and hence also by fast density-fitting techniques. This has been done starting from the VV10 Vydrov and Van Voorhis (2010a) functional, but the same can be applied to other functionals of this kind. We have implemented its density-fitting version into our molecular electronic structure code based on traditional Gaussian-type Boys (1950) functions and tested the accuracy of the approximation.
II Theory
A family of density functionals for dispersion energy correction have the form of a double space integral
$E_{6}=\int\int f({\rm\bf r}_{1},{\rm\bf r}_{2})\,\mathrm{d}^{3}{\rm\bf r}_{1}\,\mathrm{d}^{3}{\rm\bf r}_{2}$ (1)
over a function
$f({\rm\bf r}_{1},{\rm\bf r}_{2})=f\bigl(|{\rm\bf r}_{1}-{\rm\bf r}_{2}|,\rho({\rm\bf r}_{1}),\gamma({\rm\bf r}_{1}),\rho({\rm\bf r}_{2}),\gamma({\rm\bf r}_{2})\bigr)$ (2)
of the interelectronic distance $|{\rm\bf r}_{1}-{\rm\bf r}_{2}|$ and the electron density $\rho({\rm\bf r})$ and its gradient norm $\gamma({\rm\bf r})\equiv\bigl|\nabla\rho({\rm\bf r})\bigr|$ at both points ${\rm\bf r}_{1}$ and ${\rm\bf r}_{2}$. Though the function $f(r,\rho_{1},\gamma_{1},\rho_{2},\gamma_{2})$ may have a simple enough form, the integral has to be computed by six-dimensional numerical integration. We want to find a fast approximation $\tilde{E}_{6}\approx E_{6}$ as a sum of a few integrals
$\tilde{E}_{6}=\sum_{k}\int\int q_{1k}({\rm\bf r}_{1})q_{2k}({\rm\bf r}_{2})u_{k}\bigl(|{\rm\bf r}_{1}-{\rm\bf r}_{2}|\bigr)\,\mathrm{d}^{3}{\rm\bf r}_{1}\,\mathrm{d}^{2}{\rm\bf r}_{2}$ (3)
over the products of density-like distributions $q_{jk}({\rm\bf r}_{1})$ and potential-like functions $u_{k}\bigl(|{\rm\bf r}_{1}-{\rm\bf r}_{2}|\bigr)$ — opening the way to fast multipole methods or even analytical integral evaluation.
The original VV10 Vydrov and Van Voorhis (2010a) model uses a function of five variables
$f_{0}(r,\rho_{1},\gamma_{1},\rho_{2},\gamma_{2})$
$=$ $-\tfrac{3}{4}\rho_{1}\rho_{2}v_{0}\bigl(r^{2},\omega(\rho_{1},\gamma_{1}),\kappa(\rho_{1}),\omega(\rho_{2},\gamma_{2}),\kappa(\rho_{2})\bigr),$
and two spatial distributions
$\omega({\rm\bf r})$ $=$ $\omega\bigl(\rho({\rm\bf r}),\gamma({\rm\bf r})\bigr),$ (5)
$\kappa({\rm\bf r})$ $=$ $\kappa\bigl(\rho({\rm\bf r})\bigr),$ (6)
parameterized as
$\omega(\rho,\gamma)=\left(\frac{4\pi}{3}\rho+C\frac{\gamma^{4}}{\rho^{4}}\right)^{1/2},$ (7)
$\kappa(\rho)=B\rho^{1/6},$ (8)
$B\equiv\tfrac{1}{2}3^{2/3}\pi^{5/6}b$, with two adjustable parameters $C$ and $b$, and the function of the squared distance is
$v_{0}(s,\omega_{1},\kappa_{1},\omega_{2},\kappa_{2})$
$=$ $\frac{1}{(\omega_{1}s+\kappa_{1})(\omega_{2}s+\kappa_{2})\bigl((\omega_{1}+\omega_{2})s+\kappa_{1}+\kappa_{2}\bigr)}.$
Our first approximation assumes $\omega_{1}\approx\omega_{2}$, helping to simplify the function down to three variables
$f_{1}(r,\rho_{1},\gamma_{1},\rho_{2},\gamma_{2})$
$=$ $-\tfrac{3}{8}\eta(\rho_{1},\gamma_{1})\eta(\rho_{2},\gamma_{2})v\bigl(r^{2},\mu(\rho_{1},\gamma_{1}),\mu(\rho_{2},\gamma_{2})\bigr),$
with two new spatial distributions
$\eta({\rm\bf r})$ $=$ $\frac{\rho({\rm\bf r})}{\omega^{3/2}\bigl(\rho({\rm\bf r}),\gamma({\rm\bf r})\bigr)},$ (11)
$\mu({\rm\bf r})$ $=$ $\frac{\kappa\bigl(\rho({\rm\bf r})\bigr)}{\omega\bigl(\rho({\rm\bf r}),\gamma({\rm\bf r})\bigr)},$ (12)
and the new function of the distance
$v(s,\mu_{1},\mu_{2})$
$=$ $\frac{1}{(s+\mu_{1})(s+\mu_{2})\bigl(s+\tfrac{1}{2}(\mu_{1}+\mu_{2})\bigr)}.$
Besides greater simplicity, Eq. (II) allows the interpretation of $\eta({\rm\bf r})$ as a density-like quantity and leads to the geometric mean rule for the $C_{6}$ coefficients
$C_{6}=\tfrac{3}{8}\left(\int\eta({\rm\bf r})\,\mathrm{d}^{3}{\rm\bf r}\right)^{2},$ (14)
whereas the length-scale of the potential-like function $v(s,\mu_{1},\mu_{2})$ is controlled by the distribution of $\mu({\rm\bf r})$.
Our next approximation deals with the function (II)
$v(s,\mu,\nu)\approx\tilde{v}(s,\mu,\nu,\beta)$ (15)
and is of the form
$\tilde{v}(s,\mu,\nu,\beta)=\frac{1}{(s+\beta)^{3}}+\sum_{m=4}^{n}\frac{c_{m}(\mu,\nu,\beta)}{(s+\beta)^{m}}.$ (16)
At first we hoped the simplest one with only the first term might be ideal: it has the $r^{-6}$-tail the dispersion functionals are all about, and only one spatial distribution (11) is needed. This hope was forlorn after the tests on noble-gas homo- and heterodimers: the value of $\beta$ had to vary from system to system by more than twofold to reproduce well enough the potential curves. Thus the length-scale distribution (12) should be somehow accounted for.
From the asymptotic analysis
$\lim_{s\to\infty}v(s,\mu,\nu)$ $=$ $\frac{1}{s^{3}}-\frac{3(\mu+\nu)}{2s^{4}}+\dots,$ (17)
$\lim_{s\to\infty}\tilde{v}(s,\mu,\nu,\beta)$ $=$ $\frac{1}{s^{3}}+\frac{c_{4}(\mu,\nu,\beta)-3\beta}{s^{4}}+\dots,$ (18)
we get the next simplest term
$c_{4}(\mu,\nu,\beta)=3\beta-\tfrac{3}{2}(\mu+\nu)$ (19)
that may work well because it is the tail that matters.
What also matters is the sum rule, from the diagonal ($\mu=\nu$) case
$\int\limits_{0}^{\infty}r^{2}v(r^{2},\mu,\mu)\,\mathrm{d}r$ $=$ $\frac{\pi}{16\mu^{3/2}},$ (20)
$\displaystyle\int\limits_{0}^{\infty}r^{2}\tilde{v}(r^{2},\mu,\mu,\beta)\,\mathrm{d}r$ $=$ $\displaystyle\frac{\pi}{16\beta^{3/2}}+\frac{\pi c_{4}(\mu,\mu,\beta)}{32\beta^{5/2}}$ (21)
$+$ $\displaystyle\frac{5\pi c_{5}(\mu,\mu,\beta)}{256\beta^{7/2}},$
and Eq. (19) we get
$c_{5}(\mu,\mu,\beta)=\frac{16\beta^{7/2}}{5\mu^{3/2}}-8\beta^{2}+\frac{24\beta\mu}{5}.$ (22)
For $\mu\neq\nu$ we take the weighted arithmetic and geometric mean
$\begin{array}[]{lcl}c_{5}(\mu,\nu,\beta)&=&\displaystyle\frac{16\beta^{7/2}}{5(\mu\nu)^{3/4}}-8\beta^{2}\\ &+&\displaystyle(1-\zeta)\frac{24\beta(\mu\nu)^{1/2}}{5}+\zeta\frac{12\beta(\mu+\nu)}{5},\end{array}$ (23)
and by equating the derivatives
$\left.\frac{\partial^{2}\tilde{v}}{\partial\mu^{2}}\right|_{\mu=\nu=\beta}$ $=$ $\frac{3+\tfrac{6}{5}\zeta}{(s+\beta)^{5}}=\left.\frac{\partial^{2}v}{\partial^{2}\mu}\right|_{\mu=\nu=\beta},$ (24)
$\left.\frac{\partial^{2}\tilde{v}}{\partial\mu\partial\nu}\right|_{\mu=\nu=\beta}$ $=$ $\frac{3-\tfrac{6}{5}\zeta}{(s+\beta)^{5}}=\left.\frac{\partial^{2}v}{\partial\mu\partial\nu}\right|_{\mu=\nu=\beta},$ (25)
we find
$\zeta=\frac{5}{12},$ (26)
so our approximation becomes
$\begin{array}[]{lcl}\tilde{v}&=&\displaystyle\frac{1}{(s+\beta)^{3}}+\frac{\beta}{(s+\beta)^{4}}\left(3-\frac{3(\mu+\nu)}{2\beta}\right)\\ &+&\displaystyle\frac{\beta^{2}}{(s+\beta)^{5}}\left(\frac{16\beta^{3/2}}{5(\mu\nu)^{3/4}}+\frac{14(\mu\nu)^{1/2}}{5\beta}+\frac{\mu+\nu}{\beta}-8\right),\end{array}$ (27)
Figure 1 shows how it works.
Figure 1: Function $v(r^{2},\mu,\nu)$ and its approximation $\tilde{v}(r^{2},\mu,\nu,1)$ of Eq. (27) for $(\mu,\nu)=(\tfrac{1}{2},\tfrac{1}{2}),(\tfrac{1}{2},1),(1,1),(1,\tfrac{3}{2}),(\tfrac{3}{2},\tfrac{3}{2})$.
Even though $\mu({\rm\bf r})\approx\beta$ almost everywhere, it can be too big or too small somewhere, for greater numerical stability we prefer to use safer smoothed values
$\bar{\mu}({\rm\bf r})=\theta\bigl(\mu({\rm\bf r}),\beta,\alpha\bigr)$ (28)
with a filter function
$\theta(x,c,a)=c+ac\tanh\left(\frac{x-c}{ac}\right).$ (29)
Putting (27) into (II) and then (II) into (1) we get a functional of the form (3). With $\eta({\rm\bf r})$ of Eq. (11) and $\mu({\rm\bf r})$ of Eqs. (12) and (28), we need four density-like distributions
$q_{0}({\rm\bf r})$ $=$ $\eta({\rm\bf r}),$ (30)
$q_{1}({\rm\bf r})$ $=$ $\eta({\rm\bf r})\left(\beta^{-1}\mu({\rm\bf r})\right),$ (31)
$q_{2}({\rm\bf r})$ $=$ $\eta({\rm\bf r})\left(\beta^{-1}\mu({\rm\bf r})\right)^{1/2},$ (32)
$q_{3}({\rm\bf r})$ $=$ $\eta({\rm\bf r})\left(\beta^{-1}\mu({\rm\bf r})\right)^{-3/4},$ (33)
and three potential-like functions
$u_{n}(r)=\frac{\beta^{n}}{(r^{2}+\beta)^{3+n}},$ (34)
$n=0,1,2$, to compute the approximate dispersion functional
$\tilde{E}_{6}$ $=$ $\tilde{E}_{6}^{(6)}+\tilde{E}_{6}^{(8)}+\tilde{E}_{6}^{(10)},$ (35)
$\tilde{E}_{6}^{(6)}$ $=$ $-\tfrac{3}{8}U_{00}^{(0)},$ (36)
$\tilde{E}_{6}^{(8)}$ $=$ $-\tfrac{3}{8}\left(3U_{00}^{(1)}-3U_{01}^{(1)}\right),$ (37)
$\tilde{E}_{6}^{(10)}$ $=$ $-\tfrac{3}{8}\left(\tfrac{16}{5}U_{33}^{(2)}+\tfrac{14}{5}U_{22}^{(2)}+2U_{01}^{(2)}-8U_{00}^{(2)}\right),$ (38)
in terms of the integrals
$U_{ij}^{(n)}=\int\int q_{i}({\rm\bf r}_{1})u_{n}\bigl(|{\rm\bf r}_{1}-{\rm\bf r}_{2}|\bigr)q_{j}({\rm\bf r}_{2})\,\mathrm{d}^{3}{\rm\bf r}_{1}\,\mathrm{d}^{3}{\rm\bf r}_{2}.$ (39)
The numerical evaluation of the six-dimensional integrals (39) looks formally like
$U_{ij}^{(n)}={\rm\bf q}_{i}^{\mathsf{T}}{\rm\bf U}^{(n)}{\rm\bf q}_{j}.$ (40)
One way to do it is by a cubature where the values
$q_{ki}=w_{k}q_{i}({\rm\bf r}_{k}),$ (41)
$U_{kl}^{(n)}=u_{n}\bigl(|{\rm\bf r}_{k}-{\rm\bf r}_{l}|\bigr),$ (42)
are computed using the points $\{{\rm\bf r}_{k}\}$ of a three-dimensional integration grid with weights $\{w_{k}\}$, fast multipole methods Rokhlin (1985) should help here.
Another way is by density fitting
$q_{i}({\rm\bf r})=\sum_{k}b_{k}({\rm\bf r})q_{ki}$ (43)
where basis functions $\{b_{k}({\rm\bf r})\}$ are used and the coefficients are determined by the least-squares method
${\rm\bf q}_{i}={\rm\bf S}^{-1}{\rm\bf p}_{i}$ (44)
with the overlap metric
$S_{kl}=\int b_{k}({\rm\bf r})b_{l}({\rm\bf r})\,\mathrm{d}^{3}{\rm\bf r}$ (45)
and the values
$p_{li}=\int b_{l}({\rm\bf r})q_{i}({\rm\bf r})\,\mathrm{d}^{3}{\rm\bf r}$ (46)
computed by numerical integration
$p_{li}=\sum_{k}w_{k}b_{l}({\rm\bf r}_{k})q_{i}({\rm\bf r}_{k}),$ (47)
and we also calculate $q_{i}({\rm\bf r}_{k})$ from the fitted Laikov (1997) density $\rho({\rm\bf r}_{k})$ and its gradient $\gamma({\rm\bf r}_{k})$.
The matrix elements in Eq. (40) are
$U_{kl}^{(n)}=\int b_{k}({\rm\bf r}_{1})u_{n}\bigl(|{\rm\bf r}_{1}-{\rm\bf r}_{2}|\bigr)b_{l}({\rm\bf r}_{2})\,\mathrm{d}^{3}{\rm\bf r}_{1}\,\mathrm{d}^{3}{\rm\bf r}_{2}$ (48)
and can be calculated over Gaussian-type Boys (1950) basis functions in a way akin to Coulomb integrals McMurchie and Davidson (1978), we only have to learn to compute the basic integral
$\begin{array}[]{l}\displaystyle\int\int\frac{\exp\left(-a_{1}|{\rm\bf r}_{1}-{\rm\bf R}_{1}|^{2}-a_{2}|{\rm\bf r}_{2}-{\rm\bf R}_{2}|^{2}\right)}{\bigl(|{\rm\bf r}_{1}-{\rm\bf r}_{2}|^{2}+\beta\bigr)^{n+1}}\,\mathrm{d}^{3}{\rm\bf r}_{1}\,\mathrm{d}^{3}{\rm\bf r}_{2}\\ =\displaystyle\frac{\pi^{3}}{n!\,a_{1}^{3/2}a_{2}^{3/2}\beta^{n+1}}U_{0n}\left(\frac{|{\rm\bf R}_{1}-{\rm\bf R}_{2}|^{2}}{\beta},\frac{a_{1}a_{2}}{a_{1}+a_{2}}\beta\right)\end{array}$ (49)
and its partial derivatives with respect to ${\rm\bf R}_{1}$ and ${\rm\bf R}_{2}$, which can be done in terms of the special functions in two variables
$U_{mn}(x,y)=(-1)^{m}\frac{\partial^{m}}{\partial x^{m}}U_{0n}(x,y),$ (50)
$U_{mn}(x,y)=\int\limits_{0}^{\infty}\frac{y^{m+3/2}t^{m+n}}{(y+t)^{m+3/2}}\exp\left(-\frac{xyt}{y+t}-t\right)\,\mathrm{d}t.$ (51)
Our experience with the Boys Laikov (2025) functions may inspire the work on global approximations to functions (51), though for now we evaluate them by one-dimensional numerical integration using the double-exponential-like quadrature formula
$\int\limits_{0}^{\infty}p(t)\,\mathrm{d}t=\frac{1}{N}\sum\limits_{k=1}^{N-1}p\bigl(\tau(k/N)\bigr)\tau^{\prime}(k/N),$ (52)
$\tau(z)=\frac{1}{1-z}\exp\left(-\frac{a}{z}\right),$ (53)
with the parameter
$a\approx 0.56714329040978387299996866221035555$ (54)
being the solution of the equation
$a=\frac{1+a}{1+\exp(a)}.$ (55)
III Calculations
Homo- and heteroatomic noble-gas dimers are our favorite model system for testing dispersion functionals. We use our long-range-corrected version Laikov (2019a) of the PBE Perdew, Burke, and Ernzerhof (1996) exchange functional, our two-component scalar-relativistic approximation Laikov (2019b), and our L2a_3 basis set Laikov (2019c), and calculate the values shown in Table 1.
Table 1: Tests on noble-gas dimers.
system VV10 Eq. (II) Eq. (27) Eqs. (27) and (28)
$r$ $\mathcal{E}$ $\Delta r$ $\Delta\mathcal{E}$ $\Delta r$ $\Delta\mathcal{E}$ $\Delta r$ $\Delta\mathcal{E}$
HeHe 6.050 22.1 -0.007 1.2 -0.198 1.0 -0.149 1.3
HeNe 5.911 59.4 -0.012 4.8 -0.067 6.9 -0.036 8.0
NeNe 5.976 141.6 -0.013 11.5 -0.071 25.1 -0.054 25.3
HeAr 6.703 78.2 -0.014 6.6 0.028 2.7 0.005 5.1
NeAr 6.683 206.7 -0.023 22.0 -0.045 26.4 -0.042 28.6
ArAr 7.371 353.3 -0.025 28.3 -0.042 13.9 -0.035 20.0
HeKr 7.031 82.0 -0.027 9.4 0.007 5.1 -0.017 8.0
NeKr 6.952 228.5 -0.033 30.9 -0.062 36.9 -0.059 39.5
ArKr 7.593 412.5 -0.038 42.3 -0.063 27.4 -0.055 34.6
KrKr 7.789 494.5 -0.050 61.0 -0.089 47.2 -0.084 55.4
HeXe 7.522 90.4 -0.067 11.7 -0.089 3.8 -0.093 8.0
NeXe 7.324 258.9 -0.042 39.9 -0.081 41.2 -0.077 45.5
ArXe 7.922 493.5 -0.043 55.6 -0.063 31.3 -0.057 41.6
KrXe 8.099 606.9 -0.056 79.4 -0.098 55.4 -0.088 67.4
XeXe 8.384 762.5 -0.060 103.6 -0.101 67.6 -0.092 84.2
Bond lengths $r$ (bohr) and energy well depths $\mathcal{E}$ (microhartree), and their differences $\Delta r$ and $\Delta\mathcal{E}$ from the VV10 reference values.
Our first approximations of Eq. (II) works well, the errors below $1\%$ for bond lengths and below $16\%$ for bond energies are small enough. For our next working approximation of Eq. (27) we have found the sweet spot $\sqrt{\beta}=4\mbox{ bohr}$ where the errors are not much worse, mostly below $1.2\%$ ($3.3\%$ for HeHe) for bond lengths and below $18\%$ for bond energies. The further safety measures of Eq. (28) with the dimensionless $\alpha=4$ have a negligible effect. With all these approximations, we dub this new functional LPBEVVV, based on the VV10 acronym and the use of three potential-like functions.
Noble-gas dimers are the extreme example where the dispersion interaction dominates, chemically meaningful molecules and intermolecular complexes suffer much less from all these approximations.
With our density-fitting basis sets Laikov (2020) at L1a and L2a level, we have already optimized the geometries of many molecular systems and found negligible differences compared with the use of the original VV10 dispersion functional.
IV Conclusions
Our approximations to the original VV10 Vydrov and Van Voorhis (2010a) functional lead to small errors and at the same time allow much faster calculations when density-fitting is used.
It has already been (silently) applied in our study Saakian et al. (2025) of real-world organic reaction mechanisms.